16m to the 4th power - 625 =
pls help me solve this

Answer by ewatrrr(24785)


Hi,
16m^4 - 625 = 16m^4 - 5^4 = (4m^2 -25)(4m^2 + 25)=0
(4m^2 -25)(4m^2 + 25)=0
(4m^2 -25)=0
m^2 = 25/4
m = -5/2 or 5/2
(4m^2 + 25)=0
m^2 = -25/4
m = -5/2*i or 5/2*i